The cash flows from these three investments are as​ follows:

End of Year  

           A ____B ____ C
1   $1,000   $2,000   $4,000
2   2,000   2,000   4,000
3   3,000   2,000   (4,000)
4   -4,000   2,000   (4,000)
5   4,000   4,000   14,000

a. What is the present value of investment A at an annual discount rate of 11 ​percent?(Round to the nearest​ cent.)

b.  What is the present value of investment B at an annual discount rate of 11 percent?​(Round to the nearest​ cent.)

c.  What is the present value of investment C at an annual discount rate of 11 ​percent? ​(Round to the nearest​ cent.)
Answer #1

Solution-

a. Present value of investment A ( @11%) = 1000/(1.11)^1 + 2000/(1.11)^2 + 3000/(1.11)^3 + (-4000)/(1.11)^4 + 4000/1.11^5

Present value of investment A (@11%) = 1000*0.9009 + 2000*0.8116+3000*0.7312+(-4000)*0.6587 + 4000*0.5935= $4456.90

NPV of investment A =$4456.90

b. Present value of investment B (@11%) = 2000/(1.11)^1 + 2000/(1.11)^2 + 2000/(1.11)^3 + 2000/(1.11)^4 + 4000/(1.11)^5

Present value of investment B = 2000*3.1024 + 4000*0.5935 = $8578.89

NPV of investment B =$8578.89

c. Present value of investment C ( @11%) = 4000/(1.11)^1 + 4000/(1.11)^2 +(-4000)/(1.11)^3 + (-4000)/(1.11)^4 + 14000/(1.11)^5

Present value of investment C = 4000*0.9009+4000*0.8116+(-4000)*0.7312 + (-4000)*0.6587 + 14000*0.5935= $4039.80

NPV of investment C = $4039.80